How much work is necessary to compress air in an insulated cylinder from 0.20 to 0.01

m3

. Use T1= 20C and P1= 100Kpa


1
Expert's answer
2022-03-03T12:37:09-0500

Assuming the process to be adiabatic, the work can be found as follows (see https://phys.libretexts.org/Bookshelves/University_Physics/Book%3A_University_Physics_(OpenStax)/Book%3A_University_Physics_II_-_Thermodynamics_Electricity_and_Magnetism_(OpenStax)/03%3A_The_First_Law_of_Thermodynamics/3.07%3A_Adiabatic_Processes_for_an_Ideal_Gas):


"W = \dfrac{p_2V_2-p_1V_1}{1-\gamma}"

where "p_1 = 10^5Pa,V_1 = 0.2m^3,V_2=0.01m^3,\gamma=1.4" and "p_2" is the final pressure. From the adiabatic relation, the final pressure is given as follows:


"p_2 = p_1( \dfrac{V_1}{V_2})^\gamma = 10^5\cdot \dfrac{0.2^{1.4}}{0.01^{1.4}} \approx 66.29\times 10^5Pa"

Thus, obtain:


"W = \dfrac{66.29\times 10^5\cdot 0.01 - 10^5\cdot 0.2}{1-1.4} \approx -1.2\times 10^5J"

Answer. "-1.2\times 10^5J"